On a separate piece of paper (not in your lab notebook), show what volume of concentrated HCl (12 M) is needed to fully neutralize 5 mL of a 1 M solution of sodium bicarbonate (show your setup and your work). Also write the balanced chemical equation for this neutralization reaction.
Answer #1

The reaction of fully neutralization of sodium bicarbonate by HCl is as follows-

HCl(aq) + NaHCO3(aq) = NaCl(aq) + H2O(l) + CO2(g)

So one mole of HCl will completely neutralize one mole of NaHCO3.

According to the conservation of activity,

V1 × S1 = V2 × S2

[where V1=volume of HCl = ? and S1 = concentration of HCl = 12M; V2 = volume of NaHCO3 = 5mL; S2 = concentration of NaHCO3 = 1M]

So, V1 = (V2 × S2)/S1 = (5mL × 1M)/12M = 0.417mL

So volume of concentrated HCl (12M) required for fully neutralization of 5mL of 1M NaHCO3 is 0.417mL.